1. What would be the concentration (in mol/L) of a solution made by adding 200 mL of water to 50 mL of a 5.0 mol/L solution of potassium chloride?
1
Expert's answer
2021-04-27T07:48:26-0400

C1V1 = C2V2

C1 = 5.0

V1 = 50 = 0.05L

C2?

V2 200+50 = 250 = 0.25L

5 × 0.05 = 0.25 × C2

0.25 = 0.25C2

= 1 mol/L